An object of mass 0.04kg is released from rest at the lip of a smooth hemispherical bowl of radius R. The object slips down the side and collides with another object of mass 0.1 kg sitting at the bottom of the bowl. You can ignore friction between the masses and the surface of the bowl. If they stick together when they collide, how high above the bottom of the bowl will the masses go after colliding?
Answer #1

Soln mi m2 mi = 0.04 kg m₂ = 0.1kg gravity Kinetic energy of m, at bottom before collision = Work done by a migr = 0.04x 10XR=) Vs = 1.28 JR mis .: KEG of both of block = 1 (mm2) v = 5(0.0440-1) x (1.2858)² = 0.115 R J both block will .: Heighth to w